What is the equation of the hyperbola with a foci (0,-8) and (0,8) and the difference of focal radii is 10
Answer by lwsshak3(11628)

What is the equation of the hyperbola with a foci (0,-8) and (0,8) and the difference of focal radii is 10.
**
Given equation is that of a hyperbola with vertical transverse axis of the standard form:
(y-k)^2/a^2-(x-h)^2/b^2=1
For given equation:
center(0,0)
2a=given difference of focal radii=10
a=5
a^2=25
..
c=8
c^2=64
..
c^2=a^2+b^2
b^2=c^2-a^2=64-25=39
b=√39=6.24
..
Equation:
y^2/25-x^2/39=1